OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. x intercept = (6, 0); y intercept = (0, -8) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (8, 0); y intercept = (0, 6)

OpenStudy (anonymous):

First thing I do is change into the form y=mx+b.

OpenStudy (anonymous):

3y=4x-24 y=4/3x-8

OpenStudy (anonymous):

That's in y=mx+b b is the y intercept which is -8. To find the x intercept plug in y=0 in to the equation. 4x-3(0)=24 --> 4x=24 x=6

OpenStudy (anonymous):

y intercept is (0,-8) x intercept is (6,0)